Calculation: Neatly solve the question in the space provided. Demonstrate all important steps and include units. Your final answer must be written in the box to be considered correct. 1 point each. 6) A 2nC point charge and electric field form a system. If the charge moves through a potential difference of +10 V, what is the potential energy change of the system? Answer: 7) The average field between two points is 100 N/C. If the points are that are 20 cm apart, how much potential difference is there between the two points? Answer: 8) Each plate of a parallel plate capacitor has 30 pC of charge. If the capacitance of the two plates is 12 pF, what is the voltage difference between the plates? Answer: 9) How much charge is in a 8 pF capacitor if there are 50 J of potential energy stored? Answer: 10) What is the electric potential relative to infinity at a point 15cm from a +16mC point charge? Answer:
Answer #1

Solution:

Let us go to the basics first.

Question 6

Potential energy change of the system

= Vq

= 10 * (2×10^-9)

= 2*10^-8 J (Answer)

Question 7

Potential difference between the two points

= E*d

= 100 * 0.2

= 20 volts (Answer)

Question 8

Voltage difference between plates of capacitor

= Q / C

= 30*10^-6 / 12*10^-12

= 2.5*10^6 volts or 2500000 volts (Answer)

Question 9

Potential energy of capacitor ,U

= Q^2 / 2C

Thus,

Q^2

= 2CU

= 2*8*10^-12 * 50

= 8*10^-10

Thus, Q = sqrt(Q^2)

= 2.83 * 10^-5 C or 28.3 micro C (Answer)

Question 10

Electric potential is

= kq/r

= (9*10^9)*(16*10^-3)/ 0.15

= 9.6*10^8 volts (Answer)

Thanks!!!